%AL.P.Deutsch%AB.W.Lampson%TAnonlineeditor%JComm.Assoc.Comp.Mach.%V10%N 12%DDecember1967%P793-799,803%Kqed %r17%Kcstr%RComp.Sci.Tech.Rep.No.17%IBellLaboratories%CMurrayHill,NewJersey%A B.W.Kernighan%AL.L.Cherry%TASystemforTypesettingMathematics%dMay1974,revised April1977%JComm.Assoc.Comp.Mach.%Kacmcacm%V18%P151-157%DMarch1975 %TUNIXTime-SharingSystem:DocumentPreparation%Kunixbstj%AB.W.Kernighan%AM.E. Lesk%AJ.F.Ossanna%JBellSys.Tech.J.%V57%N6%P2115-2135%D1978 %AT.A.Dolotta%AJ.R.Mashey%TAnIntroductiontotheProgrammer’sWorkbench%JProc.2nd Int.Conf.onSoftwareEngineering%DOctober13-15,1976%P164-168 %TUNIXTime-SharingSystem:TheProgrammer’sWorkbench%AT.A.Dolotta%AR.C.Haight%A J.R.Mashey%JBellSys.Tech.J.%V57%N6%P2177-2200%D1978%Kunixbstj %TUNIXTime-SharingSystem:UNIXonaMicroprocessor%Kunixbstj%AH.Lycklama%JBellSys. Tech.J.%V57%N6%P2087-2101%D1978 %TTheCProgrammingLanguage%AB.W.Kernighan%AD.M.Ritchie%IPrentice-Hall%CEngle-woodCliffs,NewJersey%D1978 %TComputerRecreations%AAleph-null%JSoftwarePracticeandExperience%V1%N2%DApril-June1971%P201-204 %TUNIXTime-SharingSystem:TheUNIXShell%AS.R.Bourne%Kunixbstj%JBellSys.Tech.J. %V57%N6%P1971-1990%D1978 %AL.P.Deutsch%AB.W.Lampson%TSDS930time-sharingsystempreliminaryreferencemanual %RDoc.30.10.10,ProjectGENIE%CUniv.Cal.atBerkeley%DApril1965 %AR.J.Feiertag%AE.I.Organick%TTheMulticsinput-outputsystem%JProc.ThirdSymposium onOperatingSystemsPrinciples%DOctober18-20,1971%P35-41 %AD.G.Bobrow%AJ.D.Burchfiel%AD.L.Murphy%AR.S.Tomlinson%TTENEX,aPagedTime SharingSystemforthePDP-10%JComm.Assoc.Comp.Mach.%V15%N3%DMarch1972%Ktenex %P135-143 %AR.E.Griswold%AD.R.Hanson%TAnOverviewofSL5%JSIGPLANNotices%V12%N4%D April1977%P40-50 %AE.W.Dijkstra%TCooperatingSequentialProcesses%BProgrammingLanguages%EF.Genuys%I AcademicPress%CNewYork%D1968%P43-112 %AJ.A.Hawley%AW.B.Meyer%TMUNIX,AMultiprocessingVersionofUNIX%Kmunixunix%R M.S.Thesis%INavalPostgraduateSchool%CMonterey,Cal.%D1975 %TTheUNIXTime-SharingSystem%Kunixbstj%AD.M.Ritchie%AK.Thompson%JBellSys. Tech.J.%V57%N6%P1905-1929%D1978 %AE.I.Organick%TTheMULTICSSystem%Kmultics%IM.I.T.Press%CCambridge,Mass.%D 1972 %TUNIXforBeginners%AB.W.Kernighan%D1978 -2- %TUNIXProgrammer’sManual%AK.Thompson%AD.M.Ritchie%Kunix%IBellLaboratories%O SeventhEdition.%D1978 %AK.Thompson%TTheUNIXCommandLanguage%BStructuredProgramming—InfotechStateofthe ArtReport%IInfotechInternationalLtd.%CNicholsonHouse,Maidenhead,Berkshire,England%D March1975%P375-384%Kunix%XpwbBriefdescriptionofshellsyntaxandsemantics,withoutmuch detailonimplementation.Muchonpipesandconvenienceofhookingprogramstogether.IncludesSER-MONETTE:"Manyfamiliarcomputing‘concepts’aremissingfromUNIX.Fileshavenorecords.There arenoaccessmethods.Therearenofiletypes.Theseconceptsfillamuch-neededgap.Isincerelyhope thatwhenfuturesystemsaredesignedbymanufacturersthevalueofsomeoftheseingrainednotionsisreexamined.Likethepoliticianandhis‘commonman’,manufacturershavetheir‘averageuser’. %AJ.R.Mashey%TPWB/UNIXShellTutorial%DSeptember30,1977 %AD.F.Hartley(Ed.)%TTheCambridgeMultipleAccessSystem−UsersReferenceManual%IUniversityMathematicalLaboratory%CCambridge,England%D1968 %AP.A.Crisman(Ed.)%TTheCompatibleTime-SharingSystem%IM.I.T.Press%Kwholectssbook %CCambridge,Mass.%D1965 %TLRParsing%AA.V.Aho%AS.C.Johnson%JComp.Surveys%V6%N2%P99-124%DJune 1974 %TDeterministicParsingofAmbiguousGrammars%AA.V.Aho%AS.C.Johnson%AJ.D.Ullman %JComm.Assoc.Comp.Mach.%Kacmcacm%V18%N8%P441-452%DAugust1975 %AA.V.Aho%AJ.D.Ullman%TPrinciplesofCompilerDesign%IAddison-Wesley%CReading, Mass.%D1977 %r65%RComp.Sci.Tech.Rep.No.65%KCSTR%AS.C.Johnson%TLint,aCProgramChecker %DDecember1977%OupdatedversionTM78-1273-3%D1978 %TAPortableCompiler:TheoryandPractice%AS.C.Johnson%JProc.5thACMSymp.onPrinciples ofProgrammingLanguages%P97-104%DJanuary1978 %r39%KCSTR%RComp.Sci.Tech.Rep.No.39%IBellLaboratories%CMurrayHill,NewJersey %AM.E.Lesk%TLex—ALexicalAnalyzerGenerator%DOctober1975 %r32%KCSTR%RComp.Sci.Tech.Rep.No.32%IBellLaboratories%CMurrayHill,NewJersey %AS.C.Johnson%TYacc—YetAnotherCompiler-Compiler%DJuly1975 %TUNIXTime-SharingSystem:PortabilityofCProgramsandtheUNIXSystem%Kunixbstj%AS.C. Johnson%AD.M.Ritchie%JBellSys.Tech.J.%V57%N6%P2021-2048%D1978 %TTypingDocumentsonUNIXandGCOS:The-msMacrosforTroff%AM.E.Lesk%D1977 %AK.Thompson%AD.M.Ritchie%TUNIXProgrammer’sManual%Kunix%IBellLaboratories%O SixthEdition%DMay1975 %TTheNetworkUNIXSystem%Kunix%AG.L.Chesson%JOperatingSystemsReview%V9%N5 %P60-66%D1975%OAlsoinProc.5thSymp.onOperatingSystemsPrinciples. %TSpider—AnExperimentalDataCommunicationsSystem%Zctr127%AA.G.Fraser%JProc. IEEEConf.onCommunications%P21F%OIEEECat.No.74CH0859-9-CSCB.%DJune1974 -3- %TAVirtualChannelNetwork%AA.G.Fraser%JDatamation%P51-56%DFebruary1975 %r41%KCSTR%RComp.Sci.Tech.Rep.No.41%IBellLaboratories%CMurrayHill,NewJersey %AJ.W.Hunt%AM.D.McIlroy%TAnAlgorithmforDifferentialFileComparison%DJune1976 %AF.P.Brooks,Jr.%TTheMythicalMan-Month%IAddison-Wesley%CReading,Mass.%D1975 %XpwbReadable,classicreferenceonsoftwareengineeringandproblemsoflargeprojects,fromsomeone withexperienceinthem.Requiredreadingforanysoftwareengineer,evenifconclusionsmaynotalways beagreedwith.%br"Thesecondisthemostdangeroussystemamaneverydesigns."p.55.%br"Hence plantothrowoneaway;youwill,anyhow."p.116.%br"Cosgrovehasperceptivelypointedoutthatthe programmerdeliverssatisfactionofauserneedratherthananytangibleproduct.Andboththeactual needandtheuser’sperceptionofthatneedwillchangeasprogramsarebuilt,tested,andused."p.117. %br"Thetotalcostofmaintainingawidelyusedprogramistypically40percentormoreofthecostof developingit."p.121.%br"Asshownabove,amalgamatingproseandprogramreducesthetotalnumber ofcharacterstobestored."p.175. %TAPortableCompilerfortheLanguageC%AA.Snyder%IMaster’sThesis,M.I.T.%CCambridge, Mass.%D1974 %TTheCLanguageCallingSequence%AM.E.Lesk%AS.C.Johnson%AD.M.Ritchie%D1977 %TOptimalCodeGenerationforExpressionTrees%AA.V.Aho%AS.C.Johnson%D1975%JJ.Assoc.Comp.Mach.%Kacmjacm%V23%N3%P488-501%OAlsoinProc.ACMSymp.onTheoryof Computing,pp.207-217,1975. %AR.Sethi%AJ.D.Ullman%TTheGenerationofOptimalCodeforArithmeticExpressions%JJ.Assoc.Comp.Mach.%Kacmjacm%V17%N4%DOctober1970%P715-728%OReprintedaspp. 229-247inCompilerTechniques,ed.B.W.Pollack,Auerbach,PrincetonNJ(1972).%XpwbOptimalap-proachforstraight-line,fixednumberofregs. %TCodeGenerationforMachineswithMultiregisterOperations%AA.V.Aho%AS.C.Johnson%AJ. D.Ullman%JProc.4thACMSymp.onPrinciplesofProgrammingLanguages%P21-28%DJanuary 1977